OpenStudy (anonymous):

the number of hours per week spent drinking exeeded the number of hours per week spent studying by 1.8 hours. if an average of 18.6 hours per week were spent on both activities, determine the average number of hours spent per week spent on each activities

OpenStudy (sandra):

ok hmmm. I'm going to try to translate these into equations

OpenStudy (sandra):

so let's say time drinking is "d" and time studying is "s"

OpenStudy (sandra):

the first line means that d = s + 1.8

OpenStudy (sandra):

so if we know an average of 18.6 hours/week were spent on both activities, then that means that d+s/2 = 18.6

OpenStudy (sandra):

now you need to use substitution to get the final answer - i.e. take the value of a variable from one equation and substitute it into the other - since you need an equation to be all of one variable to solve it

OpenStudy (sandra):

in this case, how about we substitute d from the first equation (s+1.8), into the second

OpenStudy (sandra):

1. d = s + 1.8 2. (d+s)/2 = 18.6 3. (s+1.8 + s)/2 = 18.6 4. (2s + 1.8)/2 = 18.6 5. 2s + 1.8 = 37.2 6. 2s = 35.4 7. s = 17.7 17.7 hours were spent studying